Calculate the percent composition of Iron and oxygen in the product if the mass of iron is 7.874g and mass of oxygen is 1.429g


1
Expert's answer
2022-05-19T05:34:03-0400

1. We find the total mass of the product. for this we add the masses of iron and oxygen:

7.874 + 1.429 = 9.303 gr


2. We find the percentage of iron in the product. To do this, we divide the mass of iron by the mass of the total product and multiply by 100:

7.874 / 9.303 * 100 = 84.64 % Iron


3. We find the percentage of oxygen in the product. To do this, we divide the mass of oxygen by the mass of the total product and multiply by 100:

1.429 / 9.303 * 100 = 15.36 % Oxygen


ANSWER: 84.64% iron; 15.36% oxygen.